OpenStudy (anonymous):

Given f(2) = 12, which of the following could be an equation of the function? A. f(x) = 8x – 2 B. f(x) = 4x + 4 C. f(x) = 6x – 4 D. f(x) = –4x – 4

OpenStudy (anonymous):

We don't know what this is referring to, so there's no way for us to choose...

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

A. f(x) = 8x – 2 f(2) = 8(2) - 2 f(2) = 16 - 2 f(2) = 14 FALSE B. f(x) = 4x + 4 f(2) = 4(2) + 4 f(2) = 8 + 4 f(2) = 12 TRUE C. f(x) = 6x – 4 f(2) = 6(2) - 4 f(2) = 12 - 4 f(2) = 8 FALSE D. f(x) = –4x – 4 f(2) = -4(2) - 4 f(2) = -8 - 4 f(2) = -12 FALSE

OpenStudy (anonymous):

Therefore: B. f(x) = 4x + 4 f(2) = 4(2) + 4 f(2) = 8 + 4 f(2) = 12 TRUE